OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + 2x2 – 24x ?

OpenStudy (angela210793):

x3 + 2x2 – 24x=x(x2+2x-24)=x(x-4)(x+6)  so